OpenStudy (anonymous):

Factorize and explain the sollution (a-b)^2 -(c-d)^2

OpenStudy (anonymous):

cum on fast.........

OpenStudy (anonymous):

\[(a-b+c-d)*(a-b-(c-d))=(a-b+c-d)*(a-b-c+d)\]

OpenStudy (anonymous):

but (a-b)^2 = (a+b) - (a-b)

OpenStudy (anonymous):

oops (a-b)^2 = (a+b)(a-b)

OpenStudy (anonymous):

(a-b)^2=a^2-2*ab+b^2

OpenStudy (anonymous):

 Its written in my book